# Improving Alignment and Robustness with Short Circuiting

### v1

@@ 4.3 AI Agents
Short Circuiting. We mix the short circuit and retain datasets from Section 4.1 with function calling short-circuited and retain dataset. The detailed process of generating the function calling short-circuited and retain dataset is described in Section A.3. For the short-circuited LLMs, we also use the same hyperparameter configuration as in Section 4.1.
@@
Results. Figure 4 shows that after applying RR, our model is significantly more robust to harmful function calling requests, in both the no-attack and forced function-call settings, reducing harmful action compliance rates by and in the latter setting compared to baselines. Additionally, the short-circuited model retains performance on the Berkeley Function Calling Leaderboard. Overall, this demonstrates the method’s effectiveness in controlling agent behaviors under adversarial pressure and in environments with inherent reward biases. It suggests the potential for mitigating harms like power-seeking or dishonesty by short-circuiting relevant model representations, which can be as simple as adjusting the short-circuit set.
@@ 4.4 Ablation and Analysis
Ablations. We do pairwise ablations for each component of the RR loss in Table 6. First, we see that augmenting the short-circuit set with requests that bypass refusal mechanisms (w/ Augment) decreases ASR while still maintaining capabilities. Although ablating the refusal retain component (w/o refusal) increases relative robustness, it also degrades capabilities. Next, we try varying loss functions. We find that the RMU loss [29], which minimizes the distance from a constant random unit vector, fails to converge. Finally, we find that minimizing the distance from a distinct random positive vector at each step (RandP) works (though if the vector is centered at 0 (RandC), training fails). Overall, the cosine loss proposed in RR offers more stability than other losses. We then analyze the training data composition, which influences the kinds of harmful inputs the model short-circuits on.
To understand the generalization properties of short-circuiting, we split our training data into six categories of harm, train category-specific models, and measure their generalization performance across categories. We find strong in-domain generalization, indicated by the low ASR along the diagonal, and observe that training on broader categories like Harmful and Illegal Activities offers greater generalization than narrower categories like Cybercrime. We report similar ablations for Mistral-7B in Appendix F.
Representation analysis. In Figure 6, we plot the cosines between representations of the standard vs. short-circuited Llama-3-8B-Instruct model for a prefilled harmful response “Here is how to synthesize meth: 1. Start with”. We additionally plot the norms of these representations in Figure 10. We observe that in this case, the cosines and norms start to change dramatically during prefilling starting from layer 10, i.e., even before generation starts. We note that we use layers 10 and 20 for short-circuiting, so we do not expect substantial changes in the cosines and norms before layer 10 which is confirmed by the behavior of these metrics at layer 5. Although we do not directly control the representation norms during training, we observe that they often dramatically increase after short-circuiting occurs. We repeat the same experiment for Mistral-7B-Instruct and show it in Appendix G, where we also analyze two other prompts: one that leads to a similar behavior and one that triggers short-circuiting after generation starts. Importantly, we conclude that our proposed method has the intended effect on the representations and that we can detect short-circuiting by directly analyzing the internal representations. This can lead to system-level mitigations like using a probe to detect when short-circuiting happens to stop generation and, for example, provide a message that the request is considered harmful and further generation is not possible.
@@ 5 Limitations and Conclusion
Despite the promise of the methods introduced here, we emphasize that the approach we present is aiming at preventing one particular type of adversarial attack: an attack against the ability of the model to produce harmful content (often specifically against the desires of the model developer). In general, adversarial attacks can achieve other aims as well, i.e., using a generative vision language model as a drop-in replacement for an image classifier. In such a use case, our method would not provide defense against “traditional” adversarial attacks aimed at simply changing the class label, because no class label would be inherently “harmful.” Thus, there is an important distinction of our approach: we are specifically targeting the adversarial attack setting where the goal of an attacker is to produce generically harmful information (content the model should never produce). Nonetheless, for this particular use case of adversarial attacks, and for single-turn conversations that we focus on short-circuiting, our approach dramatically improves model robustness.
Overall we found that short-circuiting, based on RepE, makes models intrinsically safer and robust to unseen adversarial attacks. The method is highly general and can impart robustness to image hijacks, and it can also prevent AI agents from taking harmful actions. Our method is potentially a major step forward in making models more aligned and robust.
@@
We are thankful to Steven Basart and Stephen Casper for providing valuable feedback on the paper.
Appendix A Short Circuiting Datasets
A.1 Large Language Model Short Circuiting Dataset
To construct a dataset of diverse harmful behaviors to be short-circuited on while maintaining generalization, we prompt an uncensored LLM to generate short harmful queries and harmful completions given some examples and a wide range of categories. We then filter out all samples that have a BLEU score above 0.3 when compared to any behavior in HarmBench’s standard behaviors set [39] to avoid data contamination with the benchmark.
A.2 Multimodal Short Circuiting Dataset
To effectively construct a multimodal short-circuiting dataset containing images and their corresponding harmful queries and completions, we first use the LLaVA-Mistral-7B model [33] to generate detailed image descriptions from a sample of images from the COCO Dataset [31]. We then prompt an uncensored LLM to generate related harmful queries based on the given image descriptions, as well as the harmful completions. The final short-circuiting multimodal dataset will consist of an image and its corresponding harmful queries and harmful completions.
A.3 Function Calling Short Circuiting / Retain Dataset
To construct the Agent Short Circuiting Dataset, we start with function definitions from the Glaive Function Calling v2 [17]. Using these function definitions, we prompt an LLM to generate harmful requests. Following this, we use GPT-3.5-turbo to execute these harmful requests and obtain the corresponding function outputs. These outputs are then converted to the OpenFunctions format. Additionally, we filter out all samples that have a BLEU score above 0.1 when compared to any behavior in our proposed AgentBench (Section 4.3). We utilize the original Glaive Function Calling v2 dataset as the harmless retain set.
@@ Appendix B Refusal Evaluation
For our evaluation, we filter a subset of 500 English non-toxic user-GPT-4 requests. To measure refusal in standard models, we employ keyword checking. For the short-circuited models, we use both keyword checking and the perplexity score as measures of refusal. The refusal results are shown in Table 2. While short-circuited models show an increase in refusal rate, the rate remains considerably lower compared to more refusal-trained models like Claude-3.
@@ Appendix C Experimental Details
C.1 Additional Design Considerations for Short Circuiting
@@
Loss Multiplier Scheduling: To achieve an optimal balance between short circuiting and retention, we initially apply a large multiplier to the short circuit loss. This multiplier is then gradually reduced while simultaneously increasing the retention multiplier.
@@
Selection of Tokens for Short Circuiting: For enhanced robustness, we apply the short circuit loss to both the user and assistant text within the short circuit set for large language models and agents. For a multimodal setup, we apply short circuiting to all tokens following the image embeddings.
@@ C.2 Large Language Models
C.2.1 Detailed Short Circuiting Setup
For both models, we perform short-circuiting for 150 steps with a batch size of 16. For Mistral, we set to 5, whereas for Llama-3, we adjust to 10. Both models are trained with a batch size of 16. We specifically target layers 10 and 20 for the short-circuiting loss and insert LoRA adapters into all linear layers from layers 0 through 20. Both models are trained on 1 A100-80GB for 20 minutes.
@@ C.3 Multimodal Models
C.3.1 Detailed Short Circuiting Setup
We perform the short circuiting procedure on the language model backbone in LLaVA-NeXT-Mistral-7B [33] while freezing the image encoder and projection layer. We set to 5 and target layer 16 for the short-circuiting loss on LoRA adapters in layers 14 through 16. We alternate between short circuiting on text and multimodal training samples at each step.
@@ C.3.2 Detailed Evaluation Setup
| Physical Harm | 81.8 | 86.4 | 22.7 | 100.0 | 100.0 | 18.2 | |
| Economic Harm | 57.1 | 100.0 | 14.3 | 100.0 | 100.0 | 14.3 | |
@@ Appendix D Detailed Results in Multimodal and Agent Settings
The multimodal results on the left show that under Projected Gradient Descent (PGD) attack, short-circuiting (+s/c) is significantly more robust compared to the original model even with a safety prompt (+Prompt) that instructs the model to avoid harmful responses. Performance on multimodal capabilities benchmarks LLaVA-Wild and MMMU is preserved. In the agent setting on the right, our short-circuited model remains robust under Forced Function Calling (Forced F/C), while retaining performance on the Berkeley Function Calling Leaderboard (BFCL).
@@ Appendix F Additional Ablation Results
In general, we find that with Mistral, loss terms that leverage random vectors decrease ASR but also decrease capabilities. We don’t try the augmenting the short-circuit set or adding refusal data to the retain set because Mistral-7B is not trained with refusal mechanisms. We also find that the short-circuit set is relatively robust to distribution shifts in categories of harm.
@@ Appendix G Extended analysis of representations
Here, we provide an extended representation analysis compared to Section 4.4. In Figure 11, we use the same prefilling example (“Here is how to synthesize meth: 1. Start with”) but on the Mistral-7B models. Here we also observe that short-circuiting occurs during prefilling similarly to Llama-3-8B. However, this is not always the case: for example in Figure 12, we show a slightly modified prefilling example “Here is how to synthesize m-e-t-h: 1. Start with” (i.e., the word “meth” is purposefully obfuscated). In this case, short-circuiting happens during generation which is a common scenario: under an adversarial attack, the model starts generating harmful content but switches to meaningless outputs before harmful knowledge is revealed.
In Figure 13 and Figure 14, we show cosines and norms on Llama-3-8B and Mistral-7B for a different prefilling prompt: “Sure, here is how to make a bomb: 1. Start with”. We observe that for both model families, short-circuiting occurs during prefilling, similarly to the example presented in Section 4.4. The only difference is that the norms at layer 10 do not grow significantly, unlike the norms at layer 20 that clearly indicate short-circuiting.

### v2

@@ 4.3 AI Agents
Adding Circuit Breakers. We mix the circuit breaker and retain datasets from Section 4.1 with function calling circuit breaker and retain dataset. The detailed process of generating the function calling circuit breaker and retain dataset is described in Section A.3. For the LLMs with circuit breakers, we also use the same hyperparameter configuration as in Section 4.1.
@@
Results. Figure 4 shows that after applying RR, our model is significantly more robust to harmful function calling requests, in both the no-attack and forced function-call settings, reducing harmful action compliance rates by and in the latter setting compared to baselines. Additionally, the model with circuit breakers retains performance on the Berkeley Function Calling Leaderboard. Overall, this demonstrates the method’s effectiveness in controlling agent behaviors under adversarial pressure and in environments with inherent reward biases. It suggests the potential for mitigating harms like power-seeking or dishonesty by adding circuit breakers to the relevant model representations, which can be as simple as adjusting the circuit breaker set.
@@ 4.4 Ablation and Analysis
Ablations. We do pairwise ablations for each component of the RR loss in Table 6. First, we see that augmenting the circuit breaker set with requests that bypass refusal mechanisms (w/ Augment) decreases ASR while still maintaining capabilities. Although ablating the refusal retain component (w/o refusal) increases relative robustness, it also degrades capabilities. Next, we try varying loss functions. We find that the RMU loss [29], which minimizes the distance from a constant random unit vector, fails to converge. Finally, we find that minimizing the distance from a distinct random positive vector at each step (RandP) works (though if the vector is centered at 0 (RandC), training fails). Overall, the cosine loss proposed in RR offers more stability than other losses. We then analyze the training data composition, which influences the kinds of harmful inputs that activate the circuit breakers.
To understand the generalization properties of circuit-breaking, we split our training data into six categories of harm, train category-specific models, and measure their generalization performance across categories. We find strong in-domain generalization, indicated by the low ASR along the diagonal, and observe that training on broader categories like Harmful and Illegal Activities offers greater generalization than narrower categories like Cybercrime. We report similar ablations for Mistral-7B in Appendix F.
Representation analysis. In Figure 6, we plot the cosines between representations of the Llama-3-8B-Instruct model with and without circuit breakers for a prefilled harmful response “Here is how to synthesize meth: 1. Start with”. We additionally plot the norms of these representations in Figure 9. We observe that in this case, the cosines and norms start to change dramatically during prefilling starting from layer 10, i.e., even before generation starts. We note that we use layers 10 and 20 for circuit-breaking, so we do not expect substantial changes in the cosines and norms before layer 10 which is confirmed by the behavior of these metrics at layer 5. Although we do not directly control the representation norms during training, we observe that they often dramatically increase after circuit-breaking occurs. We repeat the same experiment for Mistral-7B-Instruct and show it in Appendix G, where we also analyze two other prompts: one that leads to a similar behavior and one that triggers circuit breakers after generation starts. Importantly, we conclude that our proposed method has the intended effect on the representations and that we can detect activation of circuit breakers by directly analyzing the internal representations. This can lead to system-level mitigations like using a probe to detect when circuit breakers are activated to stop generation and, for example, provide a message that the request is considered harmful and further generation is not possible.
@@ 5 Limitations and Conclusion
Despite the promise of the methods introduced here, we emphasize that the approach we present is aiming at preventing one particular type of adversarial attack: an attack against the ability of the model to produce harmful content (often specifically against the desires of the model developer). In general, adversarial attacks can achieve other aims as well, i.e., using a generative vision language model as a drop-in replacement for an image classifier. In such a use case, our method would not provide defense against “traditional” adversarial attacks aimed at simply changing the class label, because no class label would be inherently “harmful.” Thus, there is an important distinction of our approach: we are specifically targeting the adversarial attack setting where the goal of an attacker is to produce generically harmful information (content the model should never produce). Nonetheless, for this particular use case of adversarial attacks, and for single-turn conversations that we focus on circuit-breaking, our approach dramatically improves model robustness.
Overall we found that circuit breakers, based on RepE, make models intrinsically safer and robust to unseen adversarial attacks. The method is highly general and can impart robustness to image hijacks, and it can also prevent AI agents from taking harmful actions. Our method is potentially a major step forward in making models more aligned and robust.
@@
We are thankful to Steven Basart, Stephen Casper, David Dalrymple, and Xander Davies for providing valuable feedback on the paper.
Appendix A Circuit Breaker Datasets
A.1 Large Language Model Circuit Breaker Dataset
To construct a dataset of diverse harmful behaviors to activate circuit breakers while maintaining generalization, we prompt an uncensored LLM to generate short harmful queries and harmful completions given some examples and a wide range of categories. We then filter out all samples that have a BLEU score above 0.3 when compared to any behavior in HarmBench’s standard behaviors set [40] to avoid data contamination with the benchmark.
A.2 Multimodal Circuit Breaker Dataset
To effectively construct a multimodal circuit breaker dataset containing images and their corresponding harmful queries and completions, we first use the LLaVA-Mistral-7B model [34] to generate detailed image descriptions from a sample of images from the COCO Dataset [32]. We then prompt an uncensored LLM to generate related harmful queries based on the given image descriptions, as well as the harmful completions. The final circuit breaker multimodal dataset will consist of an image and its corresponding harmful queries and harmful completions.
A.3 Function Calling Circuit Breaker / Retain Dataset
To construct the Agent Circuit Breaker Dataset, we start with function definitions from the Glaive Function Calling v2 [17]. Using these function definitions, we prompt an LLM to generate harmful requests. Following this, we use GPT-3.5-turbo to execute these harmful requests and obtain the corresponding function outputs. These outputs are then converted to the OpenFunctions format. Additionally, we filter out all samples that have a BLEU score above 0.1 when compared to any behavior in our proposed AgentBench (Section 4.3). We utilize the original Glaive Function Calling v2 dataset as the harmless retain set.
@@ Appendix B Refusal Evaluation
For our evaluation, we filter a subset of 500 English non-toxic user-GPT-4 requests. To measure refusal in standard models, we employ keyword checking. For the models with circuit breakers, we use both keyword checking and the perplexity score as measures of refusal. The refusal results are shown in Table 2. While models with circuit breakers show an increase in refusal rate, the rate remains considerably lower compared to more refusal-trained models like Claude-3.
@@ Appendix C Experimental Details
C.1 Additional Design Considerations for Circuit Breakers
@@
Loss Multiplier Scheduling: To achieve an optimal balance between circuit-breaking and retention, we initially apply a large multiplier to the circuit-breaking loss. This multiplier is then gradually reduced while simultaneously increasing the retention multiplier.
@@
Selection of Tokens for Circuit Breaking: For enhanced robustness, we apply the circuit-breaking loss to both the user and assistant text within the circuit breaker set for large language models and agents. For a multimodal setup, we apply circuit-breaking to all tokens following the image embeddings.
@@ C.2 Large Language Models
C.2.1 Detailed Circuit Breaker Setup
For both models, we perform circuit-breaking training for 150 steps with a batch size of 16. For Mistral, we set to 5, whereas for Llama-3, we adjust to 10. Both models are trained with a batch size of 16. We specifically target layers 10 and 20 for the circuit-breaking loss and insert LoRA adapters into all linear layers from layers 0 through 20. Both models are trained on 1 A100-80GB for 20 minutes.
@@ C.3 Multimodal Models
C.3.1 Detailed Circuit Breaker Setup
We perform the circuit-breaking procedure on the language model backbone in LLaVA-NeXT-Mistral-7B [34] while freezing the image encoder and projection layer. We set to 5 and target layer 16 for the circuit-breaking loss on LoRA adapters in layers 14 through 16. We alternate between circuit-breaking on text and multimodal training samples at each step.
@@ C.3.2 Detailed Evaluation Setup
| Physical Harm | 81.8 | 86.4 | 22.7 | 100.0 | 0.0 | 18.2 | |
| Economic Harm | 57.1 | 100.0 | 14.3 | 100.0 | 0.0 | 14.3 | |
@@ Appendix D Detailed Results in Multimodal and Agent Settings
The multimodal results on the left show that under Projected Gradient Descent (PGD) attack, the model with circuit breakers is significantly more robust compared to the original model even with a safety prompt (+Prompt) that instructs the model to avoid harmful responses. Performance on multimodal capabilities benchmarks LLaVA-Wild and MMMU is preserved. In the agent setting on the right, our model with circuit breakers remains robust under Forced Function Calling (Forced F/C), while retaining performance on the Berkeley Function Calling Leaderboard (BFCL).
@@ Appendix E Multilingual Results
We do not see that trend in Table 4. We leave investigation of this to future work.
@@ Appendix F Additional Ablation Results
In general, we find that with Mistral, loss terms that leverage random vectors decrease ASR but also decrease capabilities. We don’t try the augmenting the circuit breaker set or adding refusal data to the retain set because Mistral-7B is not trained with refusal mechanisms. We also find that the circuit breaker set is relatively robust to distribution shifts in categories of harm.
@@ Appendix G Extended analysis of representations
Here, we provide an extended representation analysis compared to Section 4.4. In Figure 11, we use the same prefilling example (“Here is how to synthesize meth: 1. Start with”) but on the Mistral-7B models. Here we also observe that circuit-breaking occurs during prefilling similarly to Llama-3-8B. However, this is not always the case: for example in Figure 12, we show a slightly modified prefilling example “Here is how to synthesize m-e-t-h: 1. Start with” (i.e., the word “meth” is purposefully obfuscated). In this case, circuit-breaking happens during generation which is a common scenario: under an adversarial attack, the model starts generating harmful content but switches to meaningless outputs before harmful knowledge is revealed.
In Figure 13 and Figure 14, we show cosines and norms on Llama-3-8B and Mistral-7B for a different prefilling prompt: “Sure, here is how to make a bomb: 1. Start with”. We observe that for both model families, circuit-breaking occurs during prefilling, similarly to the example presented in Section 4.4. The only difference is that the norms at layer 10 do not grow significantly, unlike the norms at layer 20 that clearly indicate circuit-breaking.